Required information Problem 11-4A Analysis of changes in stockholders' equity accounts LO C3, P2, P3 [The following information applies to the questions displayed below.] The equity sections from Atticus Group’s 2016 and 2017 year-end balance sheets follow. Stockholders’ Equity (December 31, 2016) Common stock—$6 par value, 100,000 shares authorized, 30,000 shares issued and outstanding $ 180,000 Paid-in capital in excess of par value, common stock 140,000 Retained earnings 360,000 Total stockholders’ equity $ 680,000 Stockholders’ Equity (December 31, 2017) Common stock—$6 par value, 100,000 shares authorized, 35,200 shares issued, 4,000 shares in treasury $ 211,200 Paid-in capital in excess of par value, common stock 171,200 Retained earnings ($60,000 restricted by treasury stock) 420,000 802,400 Less cost of treasury stock (60,000 ) Total stockholders’ equity $ 742,400 The following transactions and events affected its equity during year 2017. Jan. 5 Declared a $0.50 per share cash dividend, date of record January 10. Mar. 20 Purchased treasury stock for cash. Apr. 5 Declared a $0.50 per share cash dividend, date of record April 10. July 5 Declared a $0.50 per share cash dividend, date of record July 10. July 31 Declared a 20% stock dividend when the stock’s market value was $12 per share. Aug. 14 Issued the stock dividend that was declared on July 31. Oct. 5 Declared a $0.50 per share cash dividend, date of record October 10. Problem 11-4A Part 3 3. What is the amount of the capitalization of retained earnings for the stock dividend?
Answer #1

Answer:

No. Of Shares issued & Outstanding, Dec 31, 16 = 30,000 shares

No. Of Shares purchased as Treasury Stock = 4,000 shares

No. Of Shares outstanding on the date of Stock Dividend = 30,000 - 4,000 = 26,000 shares

Stock Dividend Declared = 20%

No. Of Stocks declared as Dividend = 26,000 * 20% = 5,200 shares

Capitalization of Retained Earnings for Stock Dividend = 5,200 * $12

Capitalization of Retained Earnings for Stock Dividend = $62,400